### v2

# APOLLO: Automated LLM and Lean Collaboration for Advanced Formal Reasoning

Abstract
Formal reasoning and automated theorem proving constitute a challenging subfield of machine learning, in which machines are tasked with proving mathematical theorems using formal languages like Lean. A formal verification system can check whether a formal proof is correct or not almost instantaneously, but generating a completely correct formal proof with large language models (LLMs) remains a formidable task. The usual approach in the literature is to prompt the LLM many times (up to several thousands) until one of the generated proofs passes the verification system. In this work, we present APOLLO (Automated PrOof repair via LLM and Lean cOllaboration), a modular, model-agnostic pipeline that combines the strengths of the Lean compiler with an LLM’s reasoning abilities to achieve better proof-generation results at a low sampling budget. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub-lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top-$K$ budget. The repaired sub-proofs are recombined and reverified, iterating up to a user-controlled maximum number of attempts. On the miniF2F benchmark, we establish a new state-of-the-art accuracy of 75.0% among 7B-parameter models while keeping the sampling budget below one thousand. Moreover, Apollo raises the state-of-the-art accuracy for Goedel-Prover-SFT to 65.6% while cutting sample complexity from 25,600 to a few hundred. General-purpose models (o3-mini, o4-mini) jump from 3–7% to over 40% accuracy. Our results demonstrate that targeted, compiler-guided repair of LLM outputs yields dramatic gains in both efficiency and correctness, suggesting a general paradigm for scalable automated theorem proving.
1 Introduction
Figure 1: The summary of whole-proof generation pipeline vs. proposed Apollo agentic pipeline. LLM refers to a chosen formal theorem generator model.
Formal reasoning has emerged as one of the most challenging fields of AI with recent achievements such as AlphaProof and AlphaGeometry doing well at the International Math Olympiad competing with humans [1, 2, 3]. Formal reasoning relies both on AI models and a formal verification system that can automatically verify whether a mathematical proof is correct or not. In recent years, formal verification systems such as Lean [4] have facilitated a new form of doing mathematical research by enabling mathematicians to interact with the formal verification system and also with each other via the system, enabling larger numbers of mathematicians to collaborate with each other on a single project. As such, these formal verification systems are also called proof assistants as one can use them interactively to write a formal proof and instantaneously observe the current state of the proof and any possible errors or shortcomings in the proof generated by the compiler. Immediate access to the output of Lean compiler can help a mathematician to fix possible errors in the proof. At the same time, when a proof passes the compiler with no error, other collaborators do not need to verify the correctness of the proof. This type of collaboration is transforming the field of mathematics enabling large groups of collaborators to engage in large projects of mathematical research such as the Prime Number Theorem And More project [5]. Moreover, it has given rise to the digitization of mathematics [6].
In the AI front, large language models (LLMs) are improving at mathematical reasoning in natural language, and at the same time, they have shown a remarkable ability to learn the language of those formal verification systems and write mathematical proofs in formal language. This has led to the field of Automated Theorem Proving (ATP) where the standard approach is to prompt the LLM to generate a number of candidate proofs for a given theorem which will then be verified automatically using the formal verification system. Better models, better training sets, and better training methods has led to significant advances in this field [7, 8, 9, 10, 11].
Despite these advances, the LLMs do not really get the chance to interact with the verification system the way a human does. LLMs generate many possible proofs, sometimes as many as tens of thousands, and the verification system only marks the proofs with a binary value of correct or incorrect. This common approach is illustrated in Figure 1. When the formal verification system analyzes a proof, its compiler may generate a long list of issues including syntax errors, incorrect tactics, open goals, etc. In the current approach of the community, all of those information/feedback from the compiler remains unused, as the LLM’s proof generation is not directly engaged with the compiler errors of the formal verification system.
In this work, we address this issue by introducing Apollo, a fully automated system which uses a modular, model-agnostic pipeline that combines the strengths of the Lean compiler with the reasoning abilities of LLMs. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub-lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top-$K$ budget. The high-level overview of Apollo appears at the bottom of Figure 1.
Our contributions are as follows:
  • We introduce a novel fully automated system that directs a collaboration between LLMs, Lean compiler, and automated solvers for automated theorem proving.
  • We evaluate Apollo on miniF2F-test set using best LLMs for theorem proving and demonstrate that Apollo improves the baseline accuracies by significant margins while keeping the sampling costs much lower.
  • We establish a new SOTA on miniF2F-test benchmark for medium-sized language models. (with parameter size 7B or less)
2 Related Works
Formal theorem proving systems. At their core, formal theorem-proving (FTP) systems employ interactive theorem provers (ITPs) to verify mathematical results. In particular, Lean4 [4] is both a functional programming language and an ITP. Each theorem, lemma, conjecture, or definition must be checked by Lean’s trusted kernel, so the compiler’s verdict is binary: either the statement type-checks (True) or it fails (False). This rigorous verification dramatically increases the reliability of formal mathematics; however, it also increases the complexity of proof development: authors must both comprehend the mathematical concepts and precisely encode them in the formal language. The usefulness of Lean is also shown to go beyond theorem proving, as Jiang et al. [12] showed that Lean can be adapted to natural language logical reasoning.
Patching broken programs. Many prior works have explored using feedback to repair broken proofs. In software engineering, “repair” typically refers to program repair, i.e. fixing code that no longer runs [13]. A common source of errors is a version mismatch that introduces bugs and prevents execution. For example, SED [14] uses a neural program-generation pipeline that iteratively repairs initial generation attempts. Other notable approaches train specialized models to fix broken code based on compiler feedback (e.g. BIFI [15], TFix [16]) or leverage unsupervised learning over large bug corpora (e.g. BugLab [17]).
LLM collaboration with external experts. The use of expert feedback, whether from human users or from a proof assistant’s compiler, has proven effective for repairing broken proofs. Ringer [18] showcased automatic proof repair within the Coq proof assistant [19], enabling the correction of proofs in response to changes in underlying definitions. Jiang et al. [20] showed that leveraging automatic solvers with generative models yield better performance on formal math benchmarks. More recently, First et al. [21] showed that incorporating compiler feedback into LLM proof generation significantly improves the model’s ability to correct errors and produce valid proofs. Another line of work [22] explored an idea of mathematician and LLM collaboration, where human experts are aided by LLMs at proof writing stage.
Proof search methods. Numerous recent systems use machine learning to guide the search for formal proofs. One of the methods involves using large language models with structured search strategies, e.g. Best-First Search (BFS) [23, 24, 25] or Monte Carlo Tree Search (MCTS) [26, 27, 2]. While tree-search methods reliably steer a model toward valid proofs, they result in high inference costs and often explore many suboptimal paths before success. Another line of work leverages retrieval based systems to provide context for proof generators with potentially useful lemmas. One such example is ReProver [28] that augments Lean proof generation by retrieving relevant lemmas from a proof corpus and feeding them to an LLM, enabling the model to reuse existing proofs.
Whole proof generation. The use of standalone LLMs for theorem proving has emerged as a major research direction in automated theorem proving. One of the earliest works presented GPT-f [29], a transformer-based model for theorem proving that established the LLMs ability in formal reasoning. As training frameworks and methods advance, the community has produced numerous models that generate complete proofs without external search or tree-search algorithms. Recent work shows that both supervised models [30, 31] and those trained via reinforcement learning [30, 32, 33, 34] achieve competitive performance on formal-math benchmarks. However, whole-proof generation remains vulnerable to hallucinations that can cause compilation failures even for proofs with correct reasoning chains.
Informal Chain-of-Thought in Formal Mathematics. Several recent works demonstrate that interleaving informal “chain-of-thought” (CoT) reasoning with formal proof steps substantially improves whole-proof generation. For example, developers insert natural-language mathematical commentary between tactics, yielding large accuracy gains [31, 30, 35]. Wang, et al. [34] further show that special “thinking” tokens let the model plan its strategy and self-verify intermediate results. The “Draft, Sketch, and Prove” (DSP) [36] and LEGO-Prover [37] pipelines rely on an informal proof sketch before attempting to generate the formal proof. Collectively, these studies indicate that providing informal mathematical context, whether as comments, sketches, or dedicated tokens, guides LLMs toward correct proofs and significantly boosts proof synthesis accuracy.
3 Our Approach
Figure 2: Overview of the Apollo framework.
In this section we describe Apollo, our framework for transforming an LLM’s raw proof sketch into a fully verified Lean4 proof. Figure 2 illustrates the Apollo pipeline process with an attached Lean code before and after repair. We observe that often LLM is capable of producing a valid proof sketch; however, it often struggles with closing fine-grained goals. For example, statements h4, h5, h7, h8 are correct but the generated proofs for each sub-lemma throw compilation errors. Apollo identifies and fixes such proof steps and guides the LLM towards a correct solution.
Our compiler of choice is the Lean REPL [38], a “Read–Eval–Print–Loop” for Lean4. The REPL provides a complete list of compilation errors and warnings, each with a detailed description and the location (for example, the line number) where it occurred. Errors can include syntax mistakes, nonexistent tactics, unused tactics, and more. We aggregate this information to drive our agent’s dynamic code repair and sub-problem extraction.
3.1 Syntax Refiner
The Syntax Refiner catches and corrects superficial compilation errors: missing commas, wrong keywords (e.g. Lean3’s from vs. Lean4’s := by), misplaced brackets, etc. These syntax errors arise when a general-purpose LLM (such as o3-mini [39] or o4-mini [40]) has not been specialized on Lean4 syntax. By normalizing these common mistakes, this module ensures that subsequent stages operate on syntactically valid code.
3.2 Sorrifier
The Sorrifier patches any remaining compilation failures by inserting Lean’s sorry placeholder. We first parse the failed proof into a tree of nested proof-blocks (sub-lemmas as children). Then we compile the proof with Lean REPL [38], detect the offending line or block, and apply one of three repairs:
  1. 1.
    Line removal, if a single tactic or declaration is invalid but its surrounding block may still succeed.
  2. 2.
    Block removal, if the entire sub-proof is malformed.
  3. 3.
    Insert sorry, if the block type-checks but leaves goals open.
We repeat this procedure until the file compiles without errors. At that point, every remaining sorry marks a sub-lemma to be proved in later stages of Apollo. This part of the pipeline guarantees that the proof successfully compiles in Lean with warnings of presence of sorry statements.
Each sorry block corresponds to a sub-problem that the LLM failed to prove. Such blocks may not type-check for various reasons, most commonly LLM hallucination. The feedback obtained via REPL lets us to automatically catch these hallucinations, insert a sorry placeholder, and remove invalid proof lines.
3.3 Auto Solver
The Auto Solver targets each sorry block in turn. It first invokes the Lean4’s hint to close the goal. If goals persist, it applies built-in solvers (nlinarith, ring, simp, etc.) wrapped in try to test combinations automatically. Blocks whose goals remain open stay marked with sorry. After applying Auto Solver, a proof may be complete with no sorry’s in which case Apollo has already succeeded in fixing the proof. Otherwise, the process can repeat recursively.
3.4 Recursive repair
In the case where a proof still has some remaining sorry’s after applying the Auto Solver, Apollo can consider each of the sorry’s as a new lemma, i.e., extract its local context (hypotheses, definitions, and any lemmas proved so far), and recursively try to prove the lemmas by prompting the LLM and repeating the whole process of verification, syntax refining, sorrifying, and applying the Auto Solver.
At each of these recursive iterations, a lemma may be completely proved, or it may end up with an incomplete proof with one or a few sorry’s. This allows the Apollo to make progress in proving the original theorem by breaking down the incomplete steps further and further until the LLM or Auto Solver can close the goals. This process can continue up to a user-specified recursion depth $r$.
3.5 Proof Assembler
Finally, the Proof Assembler splices all repaired blocks back into a single file and verifies that no sorry or admit (alias for "sorry") commands remain. If the proof still fails, the entire pipeline repeats (up to a user-specified recursion depth $r$), allowing further rounds of refinement and sub-proof generation.
Apollo’s staged approach: syntax cleaning, “sorrifying,” automated solving, and targeted LLM-driven repair, yields improvements in both proof-sample efficiency and final proof correctness.
4 Experiments
In this section, we present empirical results for Apollo on the MiniF2F dataset [8], which comprises of 488 formalized problems drawn from AIME, IMO, and AMC competitions. The benchmark is evenly split into validation and test sets (244 problems each); here we report results on the MiniF2F-test partition. To demonstrate Apollo’s effectiveness, we evaluate it on a range of state-of-the-art whole-proof generation models. All experiments use Lean v4.17.0 and run on eight NVIDIA A5000 GPUs with 128 CPU cores. We used @$32$ sampling during sub-proof generation unless stated otherwise. Baseline numbers are sourced from each model’s original publication works.
4.1 The effect of applying Apollo on top of the base models
Figure 3 shows the impact of Apollo on two state-of-the-art whole-proof generation models: Goedel Prover-SFT [31] and Kimina-Prover-Preview-Distill-7B [34]. Applying Apollo increases Goedel-Prover-SFT’s top accuracy from 64.7% to 65.6% while reducing its sampling budget by two orders of magnitude (from 25,600 generated samples to only a few hundred on average). Similarly, Kimina-Prover-Preview-Distill-7B achieves a new state-of-the-art accuracy of 75.0% with roughly one-third of the previous sample budget. We report the exact numbers in Table 1.
Note that Apollo’s sampling budget is not fixed: it depends on the recursion depth $r$ and the LLM’s ability to generate sub-lemmas. For each sub-lemma that Auto Solver fails to close, we invoke the LLM with a fixed top-32 sampling budget. Because each generated sub-lemma may spawn further sub-lemmas, the total sampling overhead grows recursively, making any single @K budget difficult to report. Sub-lemmas also tend to require far fewer tactics (and thus far fewer tokens) than the main theorem. A theorem might need 100 lines of proof while a sub-lemma might need only 1 line. Hence, sampling cost does not scale one-to-one. Therefore, to compare with other whole-proof generation approaches, we report the average number of proof samples needed to reach a fully verified solution.
Figure 3: Performance of base models against the Apollo aided models on MiniF2F-test dataset at different sampling budgets ranging from 32 up to 25,600.
Table 1: Sampling cost and accuracy before vs. after applying Apollo. Cost refers to the sampling budget parameter $K$ in whole-proof generation models and average number of generated samples for Apollo. The best accuracy results are bolded.
Goedel-Prover-SFT Kimina-Prover-Preview-Distill-7B
Base Model + Apollo Base Model + Apollo
Cost Accuracy (%) Cost Accuracy (%) Cost Accuracy (%) Cost Accuracy (%)
32 57.6 32 57.6 32 63.1 32 63.1
64 59.2 80 60.7 96 68.9
3200 62.7 151 63.5 191 74.2
25,600 64.7 362 65.6 1024 70.8 307 75.0
4.2 Comparison with SoTA LLMs
Table 2 compares whole-proof generation and tree-search methods on MiniF2F-test. For each approach, we report its sampling budget, defined as the top-$K$ parameter for LLM generators or equivalent search-depth parameters for tree-search provers. Since Apollo’s budget varies with recursion depth $r$, we instead report the mean number of proof sampled during procedure.
When Apollo is applied on top of any generator, it establishes a new best accuracy for that model. For instance, Goedel-Prover-SFT’s accuracy jumps to 65.6%, surpassing the previous best result (which required 25,600 sample budget). In contrast, Apollo requires only 362 samples on average. Likewise, Kimina-Prover-Preview-Distill-7B sees a 4.2% accuracy gain while reducing its sample budget below the original 1,024. Overall, Apollo not only boosts accuracy but does so with a smaller sampling budget, setting a new state-of-the-art result among sub-7B-parameter LLMs.
We also evaluate general-purpose LLMs (OpenAI o3-mini and o4-mini). Without Apollo, these models rarely produce valid Lean4 proofs, since they default to Lean3 syntax or introduce trivial compilation errors. Yet they have a remarkable grasp of mathematical concepts, which is evident by the proof structures they often produce. By applying Apollo’s syntax corrections and solver-guided refinements, their accuracies increase from single digits (3–7%) to over 40%. These results demonstrate that in many scenarios discarding and regenerating the whole proof might be overly wasteful, and with a little help from a Lean compiler, we can achieve better accuracies while sampling less proofs from LLMs.
Table 2: Comparison of Apollo accuracy against state-of-the-art models on the miniF2F-test dataset.
Method Model size Sample budget miniF2F-test
Tree Search Methods
Hypertree Proof Search [26] 600M $64\times 5000$ $41.0\%$
InternLM2.5-StepProver + BFS + CG [24] 7B $256\times 32\times 600$ $65.9\%$
HunyuanProver v16 + BFS + DC [41] 7B $600\times 8\times 400$ $68.4\%$
BFS-Prover [25] 7B $2048\times 2\times 600$ $70.8\%$
Whole-proof Generation Methods
DeepSeek-R1-Distill-Qwen-7B [42] 7B 32 42.6%
Leanabell-Prover-GD-RL [32] 7B $128$ $61.1\%$
STP [33] 7B $25600$ $67.6\%$
o3-mini [39] - 1 $3.3\%$
$32$ $24.6\%$
o4-mini [40] - 1 $7.0\%$
Goedel-Prover-SFT [31] 7B 32 $57.6\%$
$3200$ $62.7\%$
$25600$ $64.7\%$
Kimina-Prover-Preview-Distill [34] 7B 1 $52.5\%$
$32$ $63.1\%$
$1024$ $70.8\%$
Whole-proof Generation Methods + Apollo
o3-mini + Apollo - 8 40.2% (+33.2%)
o4-mini + Apollo - 15 46.7% (+39.7%)
Goedel-Prover-SFT + Apollo 7B 362 65.6% (+0.9%)
Kimina-Prover-Preview-Distill-7B + Apollo 7B 307 75.0% (+4.2%)
4.3 Distribution of Proof Lengths
We assess how Apollo affects proof complexity by examining proof-length distributions before and after repair. Here, proof length is defined as the total number of tactics in a proof, which serves as a proxy for proof complexity.
Figure 4 presents violin plots for three base models: Kimina-Prover-Preview-Distill-7B, Goedel-Prover-SFT, and o4-mini. Each subplot shows two non-overlapping sets: proofs generated directly by the base model (“before”) and proofs produced after applying Apollo (“after”).
The proofs that Apollo succeeds in fixing in collaboration with the LLM have considerably longer proof lengths. The mean length of proofs fixed by Apollo is longer than those written by the LLM itself at least by a factor of two in each model selection scenario. This increase indicates that the proofs which the base model alone could not prove require longer, more structured reasoning chains. By decomposing challenging goals into smaller sub-lemmas, Apollo enables the generation of these extended proofs, therefore improving overall success rate.
These results demonstrate that Apollo not only raises accuracy but also systematically guides models to construct deeper, more rigorous proof structures capable of solving harder theorems.
(a) Kimina-Prover-Preview-Distill-7B
(b) Goedel-Prover-SFT
(c) OpenAI o4-mini
Figure 4: Distribution of proof lengths for selected models before (left) and after (right) applying Apollo. In the “before” plots, proof lengths cover only proofs the base model discharged without the help of Apollo; in the “after” plots, proof lengths cover only proofs completed with Apollo’s assistance.
4.4 Impact of recursion depth $r$ on proof accuracy
We evaluate how the recursion-depth parameter $r$ affects accuracy on MiniF2F-test. A recursion depth of $r=0$ corresponds to the base model’s standalone performance; higher $r$ values permit additional rounds of Apollo. Figure 5 plots accuracy versus $r$ for three models: o4-mini ($r=0\ldots 4$), Kimina-Prover-Preview-Distill-7B ($r=0\ldots 5$), and Goedel-Prover-SFT ($r=0\ldots 5$).
The o4-mini model exhibits its largest accuracy gain between $r=0$ and $r=1$, since its raw proof structure is often correct but contains many syntax errors and proof-step errors. Syntax Refiner fixes most of them and together with Auto Solver leads to a spike in accuracy. Beyond $r=1$, accuracy plateaus, since o4-mini was not explicitly designed for proving formal theorems. We note that even though accuracy gains slow down drastically for o4-mini, its accuracy still increases at a slow rate, which means that pushing parameter $r$ further has a potential to increase the efficiency of o4-mini further. In contrast, both Kimina-Prover-Preview-Distill-7B and Goedel-Prover-SFT continue to improve steadily throughout different levels of recursion depth. The results suggest that decomposing the challenging problems into smaller parts allows the model to close simpler goals, slowly building towards a complete proof.
Although the total number of proof samples grows with $r$, even $r=5$ requires only 300–400 LLM requests on average, an order of magnitude below typical budgets in the literature. Moreover, one could further increase $r$ to explore budgets in the thousands if desired.
Figure 5: Performance of Apollo on MiniF2F dataset at various recursion depth parameters $r$ on three different models: o4-mini, Kimina-Prover-Preview-Distill-7B and Goedel-Prover-SFT.
5 Limitations
Integration with tree-search methods. Our evaluation focuses exclusively on applying Apollo to whole-proof generation by LLMs and does not explore its interaction with traditional tree-search provers (e.g. BFS, MCTS). Investigating how lemma decomposition and targeted repair could accelerate or prune search paths in those systems is a promising direction for future work.
Dependence on base model’s proof sketch quality. Apollo’s ability to produce a correct formal proof largely depends on the base model and whether its initial proof has a coherent proof sketch. As [43] observe, models often tend to “cut corners,” producing proofs that are much shorter than the rigorous, multi-step arguments required to generate a valid proof. When a base model fails to write a correct proof strategy (e.g. omitting critical lemmas or suggesting irrelevant tactics, or taking a wrong approach), the Apollo is less likely to repair such a proof. Enhancing the robustness of Apollo to very poor initial sketches remains an open challenge.
6 Conclusion
In this work, we present Apollo, a novel, modular pipeline that combines syntax cleaning, automatic solvers, and LLM-driven sub-proof generation to transform an LLM’s initial proof sketch into a fully verified Lean4 proof. This framework harnesses the full power of the Lean compiler and integrated solvers, merging them with LLM systems. Applied across four whole-proof generation models, ranging from general-purpose LLMs (o3-mini, o4-mini) to specialized provers (Kimina-Prover-Preview-Distill-7B, Goedel-Prover-SFT), Apollo consistently establishes new best accuracies on the miniF2F benchmark while reducing sampling budgets by one to two orders of magnitude. Our empirical analysis shows that Apollo not only raises overall proof success rates but also guides models to generate longer, more structured proofs, as evidenced by a drastic increase in average proof length. We further demonstrate how the recursion-depth parameter $r$ trades off sample complexity against accuracy, achieving robust gains with only a few hundred samples. We believe that Apollo’s collaboration between LLMs, automatic solvers, and the Lean compiler defines a paradigm of agentic systems that produce high-quality, verifiable proofs for increasingly challenging theorems.