### v3

# APOLLO: Automated LLM and Lean Collaboration for Advanced Formal Reasoning

Abstract
Formal reasoning and automated theorem proving constitute a challenging subfield of machine learning, in which machines are tasked with proving mathematical theorems using formal languages like Lean. A formal verification system can check whether a formal proof is correct or not almost instantaneously, but generating a completely correct formal proof with large language models (LLMs) remains a formidable task. The usual approach in the literature is to prompt the LLM many times (up to several thousands) until one of the generated proofs passes the verification system. In this work, we present APOLLO (Automated PrOof repair via LLM and Lean cOllaboration), a modular, model-agnostic pipeline that combines the strengths of the Lean compiler with an LLM’s reasoning abilities to achieve better proof-generation results at a low sampling budget. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub-lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top-$K$ budget. The repaired sub-proofs are recombined and reverified, iterating up to a user-controlled maximum number of attempts. On the miniF2F benchmark, we establish a new state-of-the-art accuracy of 84.9% among sub 8B-parameter models while keeping the sampling budget below one hundred. Moreover, Apollo raises the state-of-the-art accuracy for Goedel-Prover-SFT to 65.6% while cutting sample complexity from 25,600 to a few hundred. General-purpose models (o3-mini, o4-mini) jump from 3–7% to over 40% accuracy. Our results demonstrate that targeted, compiler-guided repair of LLM outputs yields dramatic gains in both efficiency and correctness, suggesting a general paradigm for scalable automated theorem proving. The codebase is available at https://github.com/aziksh-ospanov/APOLLO
1 Introduction
Figure 1: The summary of whole-proof generation pipeline vs. proposed Apollo agentic pipeline. LLM refers to a chosen formal theorem generator model.
Formal reasoning has emerged as one of the most challenging fields of AI with recent achievements such as AlphaProof and AlphaGeometry doing well at the International Math Olympiad competing with humans [1, 2, 3]. Formal reasoning relies both on AI models and a formal verification system that can automatically verify whether a mathematical proof is correct or not. In recent years, formal verification systems such as Lean [4] have facilitated a new form of doing mathematical research by enabling mathematicians to interact with the formal verification system and also with each other via the system, enabling larger numbers of mathematicians to collaborate with each other on a single project. As such, these formal verification systems are also called proof assistants as one can use them interactively to write a formal proof and instantaneously observe the current state of the proof and any possible errors or shortcomings in the proof generated by the compiler. Immediate access to the output of Lean compiler can help a mathematician to fix possible errors in the proof. At the same time, when a proof passes the compiler with no error, other collaborators do not need to verify the correctness of the proof. This type of collaboration is transforming the field of mathematics enabling large groups of collaborators to engage in large projects of mathematical research such as the Prime Number Theorem And More project [5]. Moreover, it has given rise to the digitization of mathematics [6].
In the AI front, large language models (LLMs) are improving at mathematical reasoning in natural language, and at the same time, they have shown a remarkable ability to learn the language of those formal verification systems and write mathematical proofs in formal language. This has led to the field of Automated Theorem Proving (ATP) where the standard approach is to prompt the LLM to generate a number of candidate proofs for a given theorem which will then be verified automatically using the formal verification system. Better models, better training sets, and better training methods has led to significant advances in this field [7, 8, 9, 10, 11].
Despite these advances, the LLMs do not really get the chance to interact with the verification system the way a human does. LLMs generate many possible proofs, sometimes as many as tens of thousands, and the verification system only marks the proofs with a binary value of correct or incorrect. This common approach is illustrated in Figure 1. When the formal verification system analyzes a proof, its compiler may generate a long list of issues including syntax errors, incorrect tactics, open goals, etc. In the current approach of the community, all of those information/feedback from the compiler remains unused, as the LLM’s proof generation is not directly engaged with the compiler errors of the formal verification system.
In this work, we address this issue by introducing Apollo, a fully automated system which uses a modular, model-agnostic pipeline that combines the strengths of the Lean compiler with the reasoning abilities of LLMs. Apollo directs a fully automated process in which the LLM generates proofs for theorems, a set of agents analyze the proofs, fix the syntax errors, identify the mistakes in the proofs using Lean, isolate failing sub-lemmas, utilize automated solvers, and invoke an LLM on each remaining goal with a low top-$K$ budget. The high-level overview of Apollo appears at the bottom of Figure 1.
Our contributions are as follows:
  • We introduce a novel fully automated system that directs a collaboration between LLMs, Lean compiler, and automated solvers for automated theorem proving.
  • We evaluate Apollo on miniF2F-test set using best LLMs for theorem proving and demonstrate that Apollo improves the baseline accuracies by significant margins while keeping the sampling costs much lower.
  • We establish a new SOTA on miniF2F-test benchmark for medium-sized language models. (with parameter size 7B or less)
2 Related Works
Formal theorem proving systems. At their core, formal theorem-proving (FTP) systems employ interactive theorem provers (ITPs) to verify mathematical results. In particular, Lean4 [4] is both a functional programming language and an ITP. Each theorem, lemma, conjecture, or definition must be checked by Lean’s trusted kernel, so the compiler’s verdict is binary: either the statement type-checks (True) or it fails (False). This rigorous verification dramatically increases the reliability of formal mathematics; however, it also increases the complexity of proof development: authors must both comprehend the mathematical concepts and precisely encode them in the formal language. The usefulness of Lean is also shown to go beyond theorem proving, as Jiang et al. [12] showed that Lean can be adapted to natural language logical reasoning.
Patching broken programs. Many prior works have explored using feedback to repair broken proofs. In software engineering, “repair” typically refers to program repair, i.e. fixing code that no longer runs [13]. A common source of errors is a version mismatch that introduces bugs and prevents execution. For example, SED [14] uses a neural program-generation pipeline that iteratively repairs initial generation attempts. Other notable approaches train specialized models to fix broken code based on compiler feedback (e.g. BIFI [15], TFix [16]) or leverage unsupervised learning over large bug corpora (e.g. BugLab [17]).
LLM collaboration with external experts. The use of expert feedback, whether from human users or from a proof assistant’s compiler, has proven effective for repairing broken proofs. Ringer [18] showcased automatic proof repair within the Coq proof assistant [19], enabling the correction of proofs in response to changes in underlying definitions. Jiang et al. [20] showed that leveraging automatic solvers with generative models yield better performance on formal math benchmarks. More recently, First et al. [21] showed that incorporating compiler feedback into LLM proof generation significantly improves the model’s ability to correct errors and produce valid proofs. Another line of work [22] explored an idea of mathematician and LLM collaboration, where human experts are aided by LLMs at proof writing stage.
Proof search methods. Numerous recent systems use machine learning to guide the search for formal proofs. One of the methods involves using large language models with structured search strategies, e.g. Best-First Search (BFS) [23, 24, 25] or Monte Carlo Tree Search (MCTS) [26, 27, 2]. While tree-search methods reliably steer a model toward valid proofs, they result in high inference costs and often explore many suboptimal paths before success. Another line of work leverages retrieval based systems to provide context for proof generators with potentially useful lemmas. One such example is ReProver [28] that augments Lean proof generation by retrieving relevant lemmas from a proof corpus and feeding them to an LLM, enabling the model to reuse existing proofs.
Whole proof generation. The use of standalone LLMs for theorem proving has emerged as a major research direction in automated theorem proving. One of the earliest works presented GPT-f [29], a transformer-based model for theorem proving that established the LLMs ability in formal reasoning. As training frameworks and methods advance, the community has produced numerous models that generate complete proofs without external search or tree-search algorithms. Recent work shows that both supervised models [30, 31] and those trained via reinforcement learning [30, 32, 33, 34] achieve competitive performance on formal-math benchmarks. However, whole-proof generation remains vulnerable to hallucinations that can cause compilation failures even for proofs with correct reasoning chains.
Informal Chain-of-Thought in Formal Mathematics. Several recent works demonstrate that interleaving informal “chain-of-thought” (CoT) reasoning with formal proof steps substantially improves whole-proof generation. For example, developers insert natural-language mathematical commentary between tactics, yielding large accuracy gains [31, 30, 35]. Wang, et al. [34] further show that special “thinking” tokens let the model plan its strategy and self-verify intermediate results. The “Draft, Sketch, and Prove” (DSP) [36] and LEGO-Prover [37] pipelines rely on an informal proof sketch before attempting to generate the formal proof. Collectively, these studies indicate that providing informal mathematical context, whether as comments, sketches, or dedicated tokens, guides LLMs toward correct proofs and significantly boosts proof synthesis accuracy.
3 Our Approach
Figure 2: Overview of the Apollo framework.
In this section we describe Apollo, our framework for transforming an LLM’s raw proof sketch into a fully verified Lean4 proof. Figure 2 illustrates the Apollo pipeline process with an attached Lean code before and after repair. We observe that often LLM is capable of producing a valid proof sketch; however, it often struggles with closing fine-grained goals. For example, statements h4, h5, h7, h8 are correct but the generated proofs for each sub-lemma throw compilation errors. Apollo identifies and fixes such proof steps and guides the LLM towards a correct solution.
Our compiler of choice is the Lean REPL [38], a “Read–Eval–Print–Loop” for Lean4. The REPL provides a complete list of compilation errors and warnings, each with a detailed description and the location (for example, the line number) where it occurred. Errors can include syntax mistakes, nonexistent tactics, unused tactics, and more. We aggregate this information to drive our agent’s dynamic code repair and sub-problem extraction. The pseudo-code for our framework can be found in Algorithm 1 in the Appendix.
3.1 Syntax Refiner
The Syntax Refiner catches and corrects superficial compilation errors: missing commas, wrong keywords (e.g. Lean3’s from vs. Lean4’s := by), misplaced brackets, etc. These syntax errors arise when a general-purpose LLM (such as o3-mini [39] or o4-mini [40]) has not been specialized on Lean4 syntax. By normalizing these common mistakes, this module ensures that subsequent stages operate on syntactically valid code.
3.2 Sorrifier
The Sorrifier patches any remaining compilation failures by inserting Lean’s sorry placeholder. We first parse the failed proof into a tree of nested proof-blocks (sub-lemmas as children). Then we compile the proof with Lean REPL [38], detect the offending line or block, and apply one of three repairs:
  1. 1.
    Line removal, if a single tactic or declaration is invalid but its surrounding block may still succeed.
  2. 2.
    Block removal, if the entire sub-proof is malformed.
  3. 3.
    Insert sorry, if the block type-checks but leaves goals open.
We repeat this procedure until the file compiles without errors. At that point, every remaining sorry marks a sub-lemma to be proved in later stages of Apollo. This part of the pipeline guarantees that the proof successfully compiles in Lean with warnings of presence of sorry statements.
Each sorry block corresponds to a sub-problem that the LLM failed to prove. Such blocks may not type-check for various reasons, most commonly LLM hallucination. The feedback obtained via REPL lets us to automatically catch these hallucinations, insert a sorry placeholder, and remove invalid proof lines.
3.3 Auto Solver
The Auto Solver targets each sorry block in turn. It first invokes the Lean4’s hint to close the goal. If goals persist, it applies built-in solvers (nlinarith, ring, simp, etc.) wrapped in try to test combinations automatically. Blocks whose goals remain open stay marked with sorry. After applying Auto Solver, a proof may be complete with no sorry’s in which case Apollo has already succeeded in fixing the proof. Otherwise, the process can repeat recursively.
3.4 Recursive repair
In the case where a proof still has some remaining sorry’s after applying the Auto Solver, Apollo can consider each of the sorry’s as a new lemma, i.e., extract its local context (hypotheses, definitions, and any lemmas proved so far), and recursively try to prove the lemmas by prompting the LLM and repeating the whole process of verification, syntax refining, sorrifying, and applying the Auto Solver.
At each of these recursive iterations, a lemma may be completely proved, or it may end up with an incomplete proof with one or a few sorry’s. This allows the Apollo to make progress in proving the original theorem by breaking down the incomplete steps further and further until the LLM or Auto Solver can close the goals. This process can continue up to a user-specified recursion depth $r$.
3.5 Proof Assembler
Finally, the Proof Assembler splices all repaired blocks back into a single file and verifies that no sorry or admit (alias for "sorry") commands remain. If the proof still fails, the entire pipeline repeats (up to a user-specified recursion depth $r$), allowing further rounds of refinement and sub-proof generation.
Apollo’s staged approach: syntax cleaning, “sorrifying,” automated solving, and targeted LLM-driven repair, yields improvements in both proof-sample efficiency and final proof correctness.
4 Experiments
In this section, we present empirical results for Apollo on the MiniF2F dataset [8], which comprises of 488 formalized problems drawn from AIME, IMO, and AMC competitions. The benchmark is evenly split into validation and test sets (244 problems each); here we report results on the MiniF2F-test partition. To demonstrate Apollo’s effectiveness, we evaluate it on a range of state-of-the-art whole-proof generation models. All experiments use Lean v4.17.0 and run on eight NVIDIA A5000 GPUs with 128 CPU cores. We used @32@32 sampling during sub-proof generation unless stated otherwise. Baseline numbers are sourced from each model’s original publication works.
4.1 The effect of applying Apollo on top of the base models
Figure 3 shows the impact of Apollo on two state-of-the-art whole-proof generation models: Goedel Prover-SFT [31] and Kimina-Prover-Preview-Distill-7B [34]. Applying Apollo increases Goedel-Prover-SFT’s top accuracy from 64.7% to 65.6% while reducing its sampling budget by two orders of magnitude (from 25,600 generated samples to only a few hundred on average). Similarly, Kimina-Prover-Preview-Distill-7B achieves a new best Kimina-Prover-Preview-Distill-7B accuracy of 75.0% with roughly one-third of the previous sample budget. We report the exact numbers in Table 1.
(a) Model accuracy against sampling budget
(b) Model accuracy against generated token budget
Figure 3: Performance of base models against the Apollo aided models on MiniF2F-test dataset at different sample budgets and token length.
Note that Apollo’s sampling budget is not fixed: it depends on the recursion depth $r$ and the LLM’s ability to generate sub-lemmas. For each sub-lemma that Auto Solver fails to close, we invoke the LLM with a fixed top-32 sampling budget. Because each generated sub-lemma may spawn further sub-lemmas, the total sampling overhead grows recursively, making any single @K budget difficult to report. Sub-lemmas also tend to require far fewer tactics (and thus far fewer tokens) than the main theorem. A theorem might need 100 lines of proof while a sub-lemma might need only 1 line. Hence, sampling cost does not scale one-to-one. Therefore, to compare with other whole-proof generation approaches, we report the average number of proof samples needed to reach a fully verified solution.
Table 1: Table results comparing sampling cost, accuracy, and token usage before and after applying Apollo. “Cost” denotes the sampling budget $K$ for whole-proof generation and the recursion depth $r$ for Apollo. Column NN reports the average number of tokens generated per problem in "Chain-of-Thought" mode. Bold cells highlight the best accuracy. Results are shown on the miniF2F-test dataset for two models: Kimina-Prover-Preview-Distill-7B and Goedel-SFT.
Goedel-Prover-SFT Kimina-Prover-Preview-Distill-7B
Base Model + Apollo Base Model + Apollo
$K$ NN Acc. (%) $r$ NN Acc. (%) $K$ NN Acc. (%) $r$ NN Acc. (%)
32 16.1K 57.6% 0 16.1K 57.6% 32 140K 63.1% 0 140K 63.1%
64 31.8K 59.2% 2 38.3K 60.7% 2 406K 68.9%
3200 1.6M 62.7% 4 74.6K 63.5% 4 816K 74.2%
25600 12.7M 64.7% 6 179.0K 65.6% 1024 4.5M 70.8% 6 1.3M 75.0%
4.2 Comparison with SoTA LLMs
Table 2 compares whole-proof generation and tree-search methods on MiniF2F-test. For each approach, we report its sampling budget, defined as the top-$K$ parameter for LLM generators or equivalent search-depth parameters for tree-search provers. Since Apollo’s budget varies with recursion depth $r$, we instead report the mean number of proof sampled during procedure. Moreover, we report an average number of generated tokens as another metric for computational cost of proof generation. For some models due to inability to collect generated tokens, we leave them blank and report sampling budget only.
When Apollo is applied on top of any generator, it establishes a new best accuracy for that model. For instance, Goedel-Prover-SFT’s accuracy jumps to 65.6%, surpassing the previous best result (which required 25,600 sample budget). In contrast, Apollo requires only 362 samples on average. Likewise, Kimina-Prover-Preview-Distill-7B sees a 4.2% accuracy gain while reducing its sample budget below the original 1,024. To further validate Apollo’s generalizability, we tested the repair framework on Goedel-V2-8B [41], the current state-of-the-art theorem prover. We observe that, at similar sample and token budgets, Apollo achieves a 0.9% accuracy gain, whereas the base LLM requires twice the budget to reach the same accuracy. Overall, Apollo not only boosts accuracy but does so with a smaller sampling budget, setting a new state-of-the-art result among sub-7B-parameter LLMs.
We also evaluate general-purpose LLMs (OpenAI o3-mini and o4-mini). Without Apollo, these models rarely produce valid Lean4 proofs, since they default to Lean3 syntax or introduce trivial compilation errors. Yet they have a remarkable grasp of mathematical concepts, which is evident by the proof structures they often produce. By applying Apollo’s syntax corrections and solver-guided refinements, their accuracies increase from single digits (3–7%) to over 40%. These results demonstrate that in many scenarios discarding and regenerating the whole proof might be overly wasteful, and with a little help from a Lean compiler, we can achieve better accuracies while sampling less proofs from LLMs.
Table 2: Comparison of Apollo accuracy against state-of-the-art models on the miniF2F-test dataset. Token budget is computed over all generated tokens by LLM.
Method Model size Sample budget Token budget miniF2F-test
Tree Search Methods
Hypertree Proof Search [26] 600M $64\times 5000$ - $41.0\%$
InternLM2.5-SP+BFS+CG [24] 7B $256\times 32\times 600$ - $65.9\%$
HunyuanProver v16+BFS+DC [42] 7B $600\times 8\times 400$ - $68.4\%$
BFS-Prover [25] 7B $2048\times 2\times 600$ - $70.8\%$
Whole-proof Generation Methods
DeepSeek-R1-Distill-Qwen-7B [43] 7B 32 - 42.6%
Leanabell-Prover-GD-RL [32] 7B $128$ - $61.1\%$
STP [33] 7B $25600$ - $67.6\%$
o3-mini [39] - 1 - $3.3\%$
$32$ - $24.6\%$
o4-mini [40] - 1 - $7.0\%$
Goedel-SFT [31] 7B 32 16.1K $57.6\%$
$3200$ 1.6M $62.7\%$
$25600$ 12.7M $64.7\%$
Kimina-Prover-Preview-Distill [34] 7B 1 4.4K $52.5\%$
$32$ 140K $63.1\%$
$1024$ 4.5M $70.8\%$
Goedel-V2 [41] 8B 32 174K $83.3\%$
$64$ 349K $84.0\%$
$128$ 699K $84.9\%$
Whole-proof Generation Methods + Apollo
o3-mini + Apollo - 8 - 40.2% (+36.9%)
o4-mini + Apollo - 15 - 46.7% (+39.7%)
Goedel-Prover-SFT + Apollo 7B 362 179K 65.6% (+0.9%)
Kimina-Preview + Apollo 7B 307 1.3M 75.0% (+4.2%)
Goedel-V2 + Apollo 8B 63 344K 84.9% (+0.9%)
4.3 Distribution of Proof Lengths
We assess how Apollo affects proof complexity by examining proof-length distributions before and after repair. Here, proof length is defined as the total number of tactics in a proof, which serves as a proxy for proof complexity.
Figure 4 presents violin plots for three base models: Kimina-Prover-Preview-Distill-7B, Goedel-Prover-SFT, and o4-mini. Each subplot shows two non-overlapping sets: proofs generated directly by the base model (“before”) and proofs produced after applying Apollo (“after”).
The proofs that Apollo succeeds in fixing in collaboration with the LLM have considerably longer proof lengths. The mean length of proofs fixed by Apollo is longer than those written by the LLM itself at least by a factor of two in each model selection scenario. This increase indicates that the proofs which the base model alone could not prove require longer, more structured reasoning chains. By decomposing challenging goals into smaller sub-lemmas, Apollo enables the generation of these extended proofs, therefore improving overall success rate.
These results demonstrate that Apollo not only raises accuracy but also systematically guides models to construct deeper, more rigorous proof structures capable of solving harder theorems.
(a) Kimina-Prover-Preview-Distill-7B
(b) Goedel-Prover-SFT
(c) OpenAI o4-mini
Figure 4: Distribution of proof lengths for selected models before (left) and after (right) applying Apollo. In the “before” plots, proof lengths cover only proofs the base model proved without the help of Apollo; in the “after” plots, proof lengths cover only proofs proved with Apollo’s assistance.
Figure 5: Performance of Apollo on MiniF2F dataset at various recursion depth parameters $r$ on three different models: o4-mini, Kimina-Prover-Preview-Distill-7B and Goedel-Prover-SFT.
4.4 Impact of recursion depth $r$ on proof accuracy
We evaluate how the recursion-depth parameter $r$ affects accuracy on MiniF2F-test. A recursion depth of $r=0$ corresponds to the base model’s standalone performance; higher $r$ values permit additional rounds of Apollo. Figure 5 plots accuracy versus $r$ for three models: o4-mini ($r=0\ldots 4$), Kimina-Prover-Preview-Distill-7B ($r=0\ldots 6$), and Goedel-Prover-SFT ($r=0\ldots 6$).
The o4-mini model exhibits its largest accuracy gain between $r=0$ and $r=1$, since its raw proof structure is often correct but contains many syntax errors and proof-step errors. Syntax Refiner fixes most of them and together with Auto Solver leads to a spike in accuracy. Beyond $r=1$, accuracy plateaus, since o4-mini was not explicitly designed for proving formal theorems. We note that even though accuracy gains slow down drastically for o4-mini, its accuracy still increases at a slow rate, which means that pushing parameter $r$ further has a potential to increase the efficiency of o4-mini further. In contrast, both Kimina-Prover-Preview-Distill-7B and Goedel-Prover-SFT continue to improve steadily throughout different levels of recursion depth. The results suggest that decomposing the challenging problems into smaller parts allows the model to close simpler goals, slowly building towards a complete proof.
Although the total number of proof samples grows with $r$, even $r=5$ requires only 300–400 LLM requests on average, an order of magnitude below typical budgets in the literature. Moreover, one could further increase $r$ to explore budgets in the thousands if desired.
5 Limitations
Integration with tree-search methods. Our evaluation focuses exclusively on applying Apollo to whole-proof generation by LLMs and does not explore its interaction with traditional tree-search provers (e.g. BFS, MCTS). Investigating how lemma decomposition and targeted repair could accelerate or prune search paths in those systems is a promising direction for future work.
Dependence on base model’s proof sketch quality. Apollo’s ability to produce a correct formal proof largely depends on the base model and whether its initial proof has a coherent proof sketch. As [44] observe, models often tend to “cut corners,” producing proofs that are much shorter than the rigorous, multi-step arguments required to generate a valid proof. When a base model fails to write a correct proof strategy (e.g. omitting critical lemmas or suggesting irrelevant tactics, or taking a wrong approach), the Apollo is less likely to repair such a proof. Enhancing the robustness of Apollo to very poor initial sketches remains an open challenge.
6 Conclusion
In this work, we present Apollo, a novel, modular pipeline that combines syntax cleaning, automatic solvers, and LLM-driven sub-proof generation to transform an LLM’s initial proof sketch into a fully verified Lean4 proof. This framework harnesses the full power of the Lean compiler and integrated solvers, merging them with LLM systems. Applied across five whole-proof generation models, ranging from general-purpose LLMs (o3-mini, o4-mini) to specialized provers (Kimina-Prover-Preview-Distill-7B, Goedel-Prover-SFT, Goedel-V2), Apollo consistently establishes new best accuracies on the miniF2F benchmark while reducing sampling budgets by one to two orders of magnitude. Our empirical analysis shows that Apollo not only raises overall proof success rates but also guides models to generate longer, more structured proofs, as evidenced by a drastic increase in average successful proof length. We further demonstrate how the recursion-depth parameter $r$ trades off sample complexity against accuracy, achieving robust gains with only a few hundred samples. We believe that Apollo’s collaboration between LLMs, automatic solvers, and the Lean compiler defines a paradigm of agentic systems that produce high-quality, verifiable proofs for increasingly challenging theorems.
Appendix A Additional implementation details
Algorithm 1 presents a high-level pseudo code of the Apollo framework. We highlight that the framework is built-upon the recursive function calls up to a maximum recursion depth given by parameter $r$. $\mathbf{LLM}$ refers to a proof generator model that takes in a proof statement and outputs a formal proof attempt. $\mathbf{LeanCompiler}$ refers to Lean REPL that verifies the proof and outputs True or False. In the case of False, it also provides a detailed list of errors, their locations and types along with the error messages generated by the Lean compiler. Other functions in the pseudo code ($\mathbf{SyntaxRefiner,Sorrifier,AutoSolver,ProofAssembler}$) represent the sub-modules described in the main text. To give deeper insight into Apollo’s design, we now outline the rationale and responsibilities of each core module.
A.1 SyntaxRefiner
The SyntaxRefiner class applies a set of rule-based corrections to eliminate syntax errors in LLM-generated Lean code. During model experimentation, we collected a corpus of common mistakes, such as stray Lean 3 keywords (from, beginend), misplaced tactics, or missing delimiters, and encoded each repair as a regular-expression rule. For example:
  • Convert [from by] to [:= by].
  • Replace Lean 3 blocks (beginend) with the corresponding Lean 4 structure.
  • Ensure that commands like rw use correctly placed square brackets to avoid compilation errors.
In more complex cases, multiple regex patterns and replacements are composed to restore valid Lean 4 syntax without altering the logical content of the proof sketch.
A.2 Sorrifier
The Sorrifier module uses a feedback loop between Apollo and the Lean REPL as follows:
  1. 1.
    Apollo sends the current proof candidate to REPL.
  2. 2.
    The REPL returns a list of error messages and their source locations.
  3. 3.
    Apollo parses the proof into a syntax tree, navigates to each failing sub-lemma, and attempts to fix it or replace it with a sorry placeholder.
  4. 4.
    The updated proof is re-submitted to REPL, and this cycle repeats until no error remain.
If the REPL indicates the original theorem statement itself is malformed (e.g. due to LLM hallucination), Apollo abandons the current proof and requests a fresh generation from LLM. One example of malformed formal statement is if LLM introduces a variable in a hypothesis that was not previously defined. Such case would trigger a compilation error in the formal statement; therefore, when parsing the proof, Apollo will catch this error and request a fresh proof from LLM. To guarantee consistent variable types during lemma extraction and proof assembly, Apollo sets the following Lean options to true:
1set_option pp.instanceTypes true
2set_option pp.numericTypes true
3set_option pp.coercions.types true
4set_option pp.letVarTypes true
5set_option pp.structureInstanceTypes true
6set_option pp.instanceTypes true
7set_option pp.mvars.withType true
8set_option pp.coercions true
9set_option pp.funBinderTypes true
10set_option pp.piBinderTypes true
so that the compiler reports explicit types (e.g. $\mathbb{Z}$ vs. $\mathbb{R}$). This fail-safe prevents trivialization of theorems and ensures correct typing in the final assembled proof.
A.3 AutoSolver
The AutoSolver module takes the “sorrified” proof and applies a sequence of tactics to close as many goals as possible:
  1. 1.
    Hint-based tactics. It first invokes hint, which proposes candidate tactics. We filter out any suggestions that merely progress the goal and retain only those that fully discharge it.
  2. 2.
    Built-in tactics. If hint fails, AutoSolver systematically tries a suite of powerful Lean tactics, such as nlinarith, norm_num, norm_cast, ring_nf, simp, simp_all, and others, often combining them with try to catch failures without aborting.
  3. 3.
    Recursive fallback. Any goals that remain unsolved are reported back to Apollo. Apollo queries the REPL for the exact goal statements, spawns corresponding sub-lemmas, and then invokes the full repair loop (LLM generation, SyntaxRefiner, Sorrifier, and AutoSolver) on each sub-lemma.
This tiered strategy maximizes automation while ensuring that challenging subgoals are handled via targeted recursion.
Algorithm 1 Apollo Pseudocode for Formal Theorem Proving
1:problem statement psps, current recursion depth $r_{current}$, maximum depth $r$
2:if $r_{current}$ is not initialized then
3:  $r_{current}\leftarrow 0$
4:end if
5:if $r_{current}>r$ then
6:  return sorry $\triangleright$ If recursion depth reached, return sorry
7:end if
[…]
12:if $\mathbf{LeanCompiler}(proof)$ == True then
13:  return $\color[rgb]{0,.5,.5}proof$
14:end if
[…]
18:for $i=1$ to $n$ do
19:  $ps_{goal}\leftarrow\mathbf{GetTheoremGoal}(proof,i)$
20:  $sub\_ps\leftarrow\mathbf{TransformGoal}(ps_{goal})$
21:  $sub\_proofs[i]\leftarrow\mathbf{Apollo}(sub\_ps,r_{current},r)$
22:end for
[…]
24:return repaired_proofrepaired\_proof
Appendix B An example of Apollo proof repair procedure
Figure 6 shows a detailed run of the Apollo framework on mathd_algebra_332. First, the LLM generates an initial proof sketch, which fails to type-check in Lean. Apollo then applies its Syntax Refiner to correct simple errors (e.g. changing “from by” to “:= by”) and to strip out all comment blocks. Next, the Sorrifier replaces each failing sub-lemma with a sorry statement (six in this example). The AutoSolver module then attempts to discharge these sub-lemmas: it resolves four of them, leaving lemmas #5 and #6 unsolved. Apollo recurses on those two: each is passed again through LLM, Syntax Refiner, Sorrifier, and AutoSolver. After this single recursive iteration, all goals are closed. Finally, Apollo reassembles the full proof and returns it to the user.
This example illustrates Apollo’s repair logic. In this case the proof was fixed in one iteration; more complex theorems may require deeper recursive repair.
Figure 6: Step-by-step Apollo repair on mathd_algebra_332. The initial proof and all sub-lemma proofs are generated with o4-mini, then systematically repaired and reassembled by Apollo.
Figure 7: Proof attempt of mathd_algebra_158 produced by the base model (o4-mini) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Appendix C Comparison between base proofs generated from LLMs and Apollo-repaired counterparts
In this section, we present and discuss examples of initial proof attempts alongside their Apollo-repaired counterparts.
Figures 7, 8, 9 illustrate cases where Apollo repairs the proof without invoking the LLM again. In each example, the LLM struggles to generate the precise tactics needed to close certain goals; however, we find that regenerating the entire proof is unnecessary. Instead, built-in Lean solvers can discharge those subgoals directly. Moreover, in Figure 7, problematic block #3 is resolved simply by removing it, since the surrounding proof context is correct. Thus, omitting problematic lines can sometimes yield a valid proof.
In Figure 10, one goal is expanded via an LLM query, but the model injects extra tactics that trigger a no goals to be solved error. Apollo then repairs the first block using a combination of LLM generation and AutoSolver, while the second block is removed entirely.
Figure 11 illustrates a case where the model fails to apply nlinarith to discharge the goal. We observe that, when uncertain, the model often resorts to broad tactics in hopes of closing the goal. Here, the goal is too complex for nlinarith, so Apollo leverages both the LLM and AutoSolver to guide the proof to completion.
Figure 12 illustrates an example of proof that required a recursion depth $r=5$ to repair the initial proof attempt. We observe that LLM’s proof structure is correct; however, in many cases it over-relies on built-in solvers and rewrites. However, since the goals are too complex, this approach leads to a total of nine error blocks. Repairing those blocks requires aid from both AutoSolver module of Apollo and LLM itself. We show that if LLM grasps the correct approach, then Apollo is able to repair fine grained logic errors to produce a correct proof.
Figure 13 illustrates a case where the LLM produces an incomplete proof sketch. The first half of the proof is correct, but the model fails to discharge the remaining goal and instead uses all_goals positivity. Apollo detects this error and performs two additional recursive repair iterations to complete the proof.
Figure 14 illustrates a case where the base model takes a lazy approach by trying to close all goals with linarith. In contrast, Apollo’s repaired proof performs additional setup and discharges each goal by first proving the necessary auxiliary hypotheses. This example shows that, although the model often understands the high-level strategy, it lacks the fine-grained tactics (and compiler feedback) needed to close subgoals. Apollo decomposes the proof, identifies the failure points, and applies targeted repairs to generate a fully verified solution.
Figure 15 shows another scenario in which Apollo successfully closes the first two blocks with linarith, but the final block requires a deeper reasoning chain. Apollo augments the proof by introducing and proving the missing lemmas, which leads to a correct solution with a series of rewrites.
Figure 16 presents an example of a very large repair. As in Figure 12, Apollo requires $r=5$ to fully fix the proof. The final proof length increases from 100 lines to 216, which means Apollo applied roughly $\times 2.2$ more tactics to close the goal. The repair relied on combined effort from AutoSolver and the LLM to close all goals. We show that, even for large, complex proofs, Apollo can repair each failing sub-block and generate a correct solution.
Figure 8: Proof attempt of mathd_algebra_141 produced by the base model (o4-mini) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 9: Proof attempt of imo_1983_p6 produced by the base model (Kimina-Prover-Preview-Distill-7B) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 10: Proof attempt of induction_pord1p1on2powklt5on2 produced by the base model (Goedel-Prover-SFT) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 11: Proof attempt of mathd_algebra_293 produced by the base model (Goedel-Prover-SFT) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 12: Proof attempt of mathd_algebra_184 produced by the base model (Kimina-Prover-Preview-Distill-7B) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 13: Proof attempt of mathd_algebra_293 produced by the base model (Kimina-Prover-Preview-Distill-7B) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 14: Proof attempt of algebra_2varlineareq_fp3zeq11_3tfm1m5zeqn68_feqn10_zeq7 produced by the base model (o4-mini) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 15: Proof attempt of amc12a_2019_p12 produced by the base model (Goedel-Prover-SFT) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals.
Figure 16: Proof attempt of mathd_algebra_289 produced by the base model (Kimina-Prover-Preview-Distill-7B) versus after Apollo’s targeted repair. The left shows the original end-to-end LLM output, which does not compile in Lean; the right shows the corrected, repaired proof assembled by Apollo that closes all of the goals. Import headers are the same as other presented examples, for visibility they were omitted in this diagram.